Eliminate the arbitrary constants.

y = x + Aex + Be-2x

1
Expert's answer
2021-12-13T05:57:05-0500

"Given\mathrm{:}\ \ y=x+Ae^x+Be^{-\mathrm{2}x}....................\left(\mathrm{1}\right) \\

We\ \ need\ \ to\ \ e\mathrm{lim}inate\ \ the\ \ two\ \ cons\mathrm{tan}ts\ \ resulting\ \ \ \ from \\

a\ \ \mathrm{sec}ond\ \ order\ \ ode. \\

y'=\mathrm{1}+Ae^x-\mathrm{2}Be^{-\mathrm{2}x}\ \ \ \ \ ..............\left(\mathrm{2}\right) \\

 \\

y''=Ae^x+\mathrm{4}Be^{-\mathrm{2}x}......................\left(\mathrm{3}\right) \\

 \\

add\ \ \left(\mathrm{2}\right)\ \ and\ \ \left(\mathrm{3}\right)\ \ \\

 \\

we\ \ have\ \ y''+y'=\mathrm{1}\ +\mathrm{2}\ Ae^x+\mathrm{2}Be^{-\mathrm{2}x}\ \ .............\left(\mathrm{4}\right) \\

 \\

multiply\ \ \left(\mathrm{1}\right)\ \ by\ \ \mathrm{2,} \\

 \\

we\ \ \ have\ \ \ \mathrm{2}y=\mathrm{2}x+\mathrm{2}Ae^x+\mathrm{2}Be^{-\mathrm{2}x}...................\left(\mathrm{5}\right) \\

 \\

subtract\ \ \ \left(\mathrm{5}\right)\ \ \ from\ \ \left(\mathrm{4}\right) \\

 \\

y''+y'-\mathrm{2}y=\mathrm{1}\ +\mathrm{2}\ Ae^x+\mathrm{2}Be^{-\mathrm{2}x}\ -\mathrm{2}\ Ae^x-\mathrm{2}Be^{-\mathrm{2}x}-\mathrm{2}x \\

 \\

y''+y'-\mathrm{2}y=\mathrm{1}\ -\mathrm{2}x \\"